Home‎ > ‎Reports‎ > ‎Crew100A‎ > ‎

day03


Crew 100A Log Book for February 14, 2011

Photos of the day





Commander's Check-In Report

Ivo Ferreira Reporting

Crew Physical Status: Good

Time Departed/Returned from EVA #2: 17:27 pm / 18:00 pm

Brief Narrative of Field Mission Results: We conducted a series of habitability experiments during the morning.

The rover was tested outside at a small track made for the occasion (an analog of the F1 Cataluña Grand Prix). The first tests were done to study the influence of training on the performance by conducting test with experts and non-experts in the rover operation. Two of the crew members conducted the first of a series of 8 EVAs to study the stress related to communication during an EVA. Their heart rate was being monitored to measure the effect of communication in their stress/physical effort. The mission was a total success.

EVA Data/Interpretations: N/A (To be completed in the following days)

Engineering/Hab Maintenance: N/A

Plans for Tomorrow: Continue with experiments: habitability, rover and drone and EVA to test stress level in the afternoon. Observatory training.

Reports scheduled for today: Engineering, Commander, Journalist and EVA#2

Inventory: None

Support Requested: N/A

Miscellaneous: Today with had our first contact with the Martian species :- ) We got an unexpected visit from local motorbike riders very eager to know about the MDRS. We tour them around and explained some of the experiments and they were really surprised with the Habitat. Nothing like spreading some interest in space related subjects. On an unrelated subject we received several kitchen items and food supplies sent ahead for crew 101.



Commander's Report

Ivo Ferreira Reporting

Today we started the "real" research.

We started the day with Kévin and Jeffrey handling the equipment they slept with to another two crew members. Tonight it will be Audrey and Romain who will have the chance to share their night with a Holter monitor, a temperature recorder and an experimental device to

non-invasively determine parameters such as heart rate, cardiac output, stroke volume, etc. They will be their Valentine for the night.

We shared our saliva with the science community and went on to the first of a series of habitability experiments run by Ayako. In complete isolation we had to smell different fragrances and measure our reaction both subjectively and objectively (more saliva).

The afternoon was reserved to the first teleoperation experiment with Romain as the non-expert user, and Jeffrey as the "Rover-master". There were some differences but both of them managed to complete our We are well below the record on the real Cataluña track… they should update their records database...

mini-F1 analog track (set up for the occasion) in less than 3 minutes.

Later in the afternoon, we conducted the first of a set of 8 EVAs to study the effect of the stress added by communication during an EVA. Kévin and Audrey, conducted a difficult EVA with a lot of constraints with me acting as the Mission Control. They did quite well and managed to save all the crew for certain (simulated) death!

Tomorrow there will be drones flying in the Desert. UFO sightings awaited!


Engineering Report

Kévin Theatre Reporting

Generator/Electricity:

· Kitty (Diesel Generator) (Used? Yes/No): Yes

· Honey (Gas Generator) (Used? Yes/No): No

· Battery Charge Level: Checked by DG

o 10.59 V everstart

o 12.37 V armor plate

o 13.32 V green


Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off):

1 generator run, all is good


Fuel Status:

· Diesel (Full/0.75/0.5/0.25/Empty): 0.85

· Propane Tank (Full/0.75/0.5/0.25/Empty): 0.61

· Gasoline Tank (Full/0.75/0.5/0.25/Empty): 12 Gallons

Notes/Comments:


Oil:

· Kitty Oil Quantity (# of Quarts in storage): Leave it to DG

· Honey Oil Quantity (# of Quarts in storage): Leave it to DG

· ATV Oil Quantity (# of Quarts in storage): 6 quarts on the engineering work bench

· Car Oil Quantity (# of Quarts in storage):


Water Status:

· Outside Potable Water Tank Level (inches from bottom): almost Full

· Trailer Potable Water Tank Level (inches from bottom): 1/3 Full

· Hab Potable Water Tank Level (inches from bottom): 9

· Potable Water Meter Reading: 420479


Notes/Comments


GreenHab:

· Condition of plants in Tank 1: green water with froth (bubbles)

· Condition of plants in Tank 2: green things

· Condition of plants in Tank 3: green things

· Greenhab notes/ comments:


Observatory:

· Telescope Used? (Yes/no) No

Observatory Notes:


Transportation:

Hab Car (NOT RENTAL)

· Used (yes/no): No

· Oil Checked (yes/no): No

· Tire Status: Not checked

· Odometer at end of day: 123611.6

· Notes/Comments on Hab Car:


ATV 1 (Opportunity):

· Used: No

· Oil Checked: no

· Fuel Consumed: Replaced

· Tire Status: Filled

· Comments:


ATV 2 (Spirit):

· Used: No

· Oil Checked: No

· Fuel Consumed:

· Tire Status:

· Comments: Still not fixed


ATV 3 (Viking I):

· Used: No

· Oil Checked: no

· Fuel Consumed: Replaced

· Tire Status: Filled

· Comments:


ATV 4 (Viking II):

· Used: No

· Oil Checked: No

· Fuel Consumed: Replaced

· Tire Status: Filled

· Comments:


Heating and Ventilation:

· Thermostat setting upstairs (Farenheit): 65

· Actual temperature upstairs (Farenheit): 75

· Thermostat setting downstairs (Farenheit): 55

· Actual temperature downstairs (Farenheit): 64


Computers/Networking Infrastructure:

Notes/comments on internet/computers?: All is well

General Engineering

1. Notes/Comments:


EVA 2 Report

Kévin Theatre Reporting

EVA #2, 17:30- 18:30

Crew Members: Ivo Ferreira, Audrey Lan San, Kévin Théatre

Site Location: Foot: 518135 4250545,518087 4250490,0518153 4250453

Transit Mode: Foot

Objectives: Mission 1 (testing the influence of stress related to communication on an EVA).

Results: Managed to comply with the 2 critical objectives. 1 scientific objective failed.

Lessons Learned: Don't hide too much a so-called sample, else your team won't find it. Good radios are required. It's really difficult to make you understood otherwise.


Chef's Report

Ivo Ferreira Reporting

1a) Today was: Non-cooking day

1b) Was today a special day celebrated at the main meal with special food or activity? No.

2)  List the foods served at today's main meal (usually the evening meal), giving full names of each. Chicken Kung Fu, Tomato Powder, Dried onions, Cookies,No special beverage.

3) List any main meal foods not finished by the crew (leftovers stored for later use or discarded)

Example:  Nothing

4)  Recipes for anything more complicated than following package directions exactly, or rehydrating.  Everything followed the instructions on the package.

5)  Lessons learned  N/A

6) Comments/questions for Kim and Jean, the food study investigators: What is the different between non cooking and pre prepared food?


Journalist Report

Romain Benchenafi Reporting

Third day under the sunny sky of Mars for the crew 100A. We designed a track in order to experiment the Rover's maneuverability, its sensibility to obstacles and the telecommunication interface. A second EVA has been

performed aiming at studying communication related to stress. Time is coming to analyse the results of our experiment and to develop the dynamic of our team. Hope next day will be as good as this one!